OpenStudy (anonymous):

is 16,8,4,2 arithmetic or geometric?

hartnn (hartnn):

what do you think ? you get common ratio or common difference between the terms?

OpenStudy (anonymous):

i think it would be geometric.... am i correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what about 2,3,5,8,12? is it arithmetic or geometric? explain.

hartnn (hartnn):

geometric is correct :) how did u figure it out?

OpenStudy (anonymous):

i think its arithmetic but im not sure why

OpenStudy (anonymous):

ugh i dont feel like explaining

hartnn (hartnn):

for arithmetic, there's a common difference between the terms

hartnn (hartnn):

like d =a2-a1 = a3 -a2 = a4-a3 = ....= and so on

OpenStudy (anonymous):

so add, 1,2,3,4... etc? so it WOULD be arithmetic

hartnn (hartnn):

in case of 2,3,5,8,12 are the differences SAME? ?

OpenStudy (anonymous):

no?

hartnn (hartnn):

does 3-2 equal 5-3 equal 8-5

hartnn (hartnn):

yes, 'no' is correct, so its not arithmetic

OpenStudy (anonymous):

so its geo?

hartnn (hartnn):

for geometric there is a common ratio, r = a2/a1 = a3/a2 = ....and so on

hartnn (hartnn):

so does 3/2 equal 5/3 ?

OpenStudy (anonymous):

ok ur making me confused.

hartnn (hartnn):

instead of taking differences (for arithmetic), you just take ratios(for geometric)!

OpenStudy (anonymous):

so its geometric again eh?

hartnn (hartnn):

for 2,3,5,8 .... you don't get common difference you don't get common ratio too!! so its neither :)

OpenStudy (anonymous):

2,3,5,8,12 you mean?

OpenStudy (anonymous):

ok thank hartnn

hartnn (hartnn):

yes. and you can look here for some explanation on this topic :) Tutorial on AP,GP,HP. http://openstudy.com/study#/updates/503bb2a0e4b007f9003103b0

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

thanks for help goku

hartnn (hartnn):

goku is always happy to help :D

OpenStudy (anonymous):

TODAYS EPISODE GOKU SUMMON THE WATERMELON XD

OpenStudy (anonymous):

wanna help me with another question?

OpenStudy (anonymous):

nevermind

hartnn (hartnn):

ask it in a new post :) if i am not available, then others can help!

OpenStudy (anonymous):

you good at science?

OpenStudy (anonymous):

ima switch to science